please can somebody show me how to solve this
I need to find 3 solutions
x^3+9x^2+14x=0
please show workings so I can understand

Answer by htmentor(1343)

x^3+9x^2+14x=0
The equation can be factored as
x(x^2 + 9x + 14) = 0
The expression inside the parentheses can be further factored as
(x+7)(x+2)
Writing the fully-factored equation we can immediately read off the 3 solutions:
x(x+7)(x+2) = 0
So the solutions are x = 0, x = -7 and x = -2